(no subject)

Emmanuel Monnier (monnier@hep.uchicago.edu)
Wed, 22 Sep 1999 15:41:28 -0500 (CDT)

Shift Summary for owl shift, 09/22/99

On Shift: E. Monnier, J. Ladue

Hours of beam available at all intensities: 8
Hours of beam available at nominal intensity: 5.3
Hours of E799 data (special runs not included): 5.3
Hours of special runs (laser scans, muon runs...): 0.5

Accounting of lost time:

spurious beam, not very stable.

========================================================================
Runs ended this shift:
Run 14665 - 13.7 M events - BEAM_799

Tapes brought to FCC this shift:
(including tapes for online split)
10 tapes, UPA-C:16-18 and UPS007

========================================================================
Accesses during shift:

none

========================================================================
Overview of shift:
E799 datataking with really lousy, intermittent beam, DAQ_DEAD...

========================================================================

9:40 BoBT request a change to high intensity w 2 turns (9E12)

11:00 BoA up to 1.9 call mcr and complain
11:20 Still no improvement
11:36 BoA now 1.2 spill 246
11:54 BoA bort intermittant beam to users very bad boa and SEM <4e12>
13:20 Good sem <8e12> and boa 1.2, spill 296
15:03 we have mounted during the shift the following split tapes
UPM004 UPF003 UPT001 UPH001 UPX002

========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift

NM2SR is set at 2650 Amps, readback is only 2570 A.

Special run, need to do a run for 30' with RC threshold at 100 MeV

Tapes need to be balanced.